# Stable Offline Value Function Learning with Bisimulation-based Representations

### v3

@@ C.2 Baselines
It is a bisimulation-based metric learning algorithm that is off-policy evaluation variant of mico (Castro et al., 2022; Pavse & Hanna, 2023a). It learns representations directly such that the $\pi$-bisimulation metric is satisfied and makes no assumptions on the transition dynamics. Its additional learning rate is the output dimension of $\phi$.
@@
The dr3 regularizer minimizes the total feature co-adaptation by adding the term $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$ as an auxiliary task to the main fqe loss (Kumar et al., 2021). (Ma et al., 2024) introduced an improvement to this auxiliary loss by suggesting that the absolute value of the feature co-adaptation be minimized, i.e., $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}|\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})|$. We use $\alpha=0.1$ as its auxiliary task weight. Absolute dr3’s only hyperparameters are the auxiliary task weight $\alpha$ and the $\phi$ output dimension.
@@ C.3 Additional Results
C.3.1 Understanding the Stability of the krope Learning Procedure
In this section analyze how the deadly triad affects krope learning procedure. Since krope’s learning algorithm is a semi-gradient method, it is susceptible to instability (Tsitsiklis & Van Roy, 1997). As such, these experiments aim to understand when we might expect krope to diverge and when it might mitigate the divergence of fqe.
We conduct our experiments on the Markov reward process in Figure 5(a), which was introduced by (Feng et al., 2019). The mrp consists of $4$ non-terminal states, $1$ terminal state (the box), and only $1$ action. The value function estimate is linear in the weight $w=[w_{1},w_{2},w_{3}]$, so the native features of the states are $[1,0,0]$, $[0,1,0]$ , $[0,0,2]$, and $[0,0,1]$ (clockwise starting from left-most). We set $\gamma=1$, so the optimal weights are $w^{*}=[0.8,1,0]$. In this setup, we say a transition is a bad transition if the bootstrapping target is a moving target for the current state (Asadi et al., 2024). For example, the transition from $w_{3}$ to $2w_{3}$ is a bad transition since updates made to $w_{3}$ may move $2w_{3}$ further away. When this transition is sampled at a frequency that is different from the on-policy distribution, algorithms such as td, lspe, and fqe tend to diverge (Asadi et al., 2024).
(a) Divergence Counterexample
(b) krope Training loss
(c) fqe+krope Training loss
Figure 5: Figure 5(a): Markov reward process counterexample designed to illustrate divergence; $r$ denotes the rewards and $p$ denotes the probability of transition (Feng et al., 2019). Figures 5(b) and 5(c): krope training loss and fqe+krope training loss vs. epochs respectively when different datasets are paired with $\mathcal{D}_{1}$; results are averaged over $20$ trials, shaded region is the $95\%$ confidence interval, and lower is better.
To better understand the stability of the learning procedures, we design the following experiment. We start with two datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. $\mathcal{D}_{1}$ consists of $2000$ on-policy transitions and $5000$ off-policy bad transitions from $w_{3}$ to $2w_{3}$, while $\mathcal{D}_{2}$ consists of only $2000$ on-policy transitions. We then consider four different variations of $\mathcal{D}_{2}$: $\mathcal{D}_{2}^{w_{1}}$, $\mathcal{D}_{2}^{w_{2}}$, $\mathcal{D}_{2}^{w_{3}}$, and $\mathcal{D}_{2}^{2w_{3}}$, where each variation denotes a dataset where we add $5000$ off-policy transitions from the specified state to $\mathcal{D}_{2}$. With this setup, our goal is understand which dataset pairing with $\mathcal{D}_{1}$: $(\mathcal{D}_{1},\mathcal{D}_{2}^{w_{1}})$, $(\mathcal{D}_{1},\mathcal{D}_{2}^{w_{2}})$, $(\mathcal{D}_{1},\mathcal{D}_{2}^{w_{3}})$, and $(\mathcal{D}_{1},\mathcal{D}_{2}^{2w_{3}})$ increases the susceptibility of krope to divergence. Recall that at each training step, krope samples pairs of transitions. Thus, in this setup, krope will sample one transition from $\mathcal{D}_{1}$ and the other transition from the other dataset.
In Figures 5(b) and 5(c), we show the training loss of krope only and fqe+krope with krope as an auxiliary task. In both cases, the sampled state-actions are fed into a linear network encoder with a bias component and no activation function which outputs a $d=3$ representation. In case of fqe, this representation is then fed into a linear function to output the scalar value. From Figure 5(b), we find that even though $\mathcal{D}_{1}$ consists of mostly bad transitions, if krope also samples from $\mathcal{D}_{2}^{w_{1}}$ or $\mathcal{D}_{2}^{w_{2}}$, its divergence is mitigated. However, as expected, if krope uses $\mathcal{D}_{2}^{w_{3}}$ or $\mathcal{D}_{2}^{2w_{3}}$, krope diverges since the pairing of samples from $(\mathcal{D}_{1},\mathcal{D}_{2}^{w_{3}})$, and $(\mathcal{D}_{1},\mathcal{D}_{2}^{2w_{3}})$ leads to krope chasing a moving bootstrapped target, resulting in divergence. Building upon this insight, we can then better understand when krope can mitigate fqe’s divergence. From Figure 5(c), we see the expected result that when fqe uses $\mathcal{D}_{1}$ only, it diverges since $\mathcal{D}_{1}$ consists of many bad transitions. However, if fqe uses krope as an auxiliary loss, krope may either reduce the divergence $(\mathcal{D}_{2}^{w_{1}}\text{ and }\mathcal{D}_{2}^{w_{2}})$ or worsen it $(\mathcal{D}_{2}^{w_{3}}\text{ and }\mathcal{D}_{2}^{2w_{3}})$ depending on which dataset is paired with $\mathcal{D}_{1}$. While in practice, it is unclear to determine which situation is more likely, this result sheds light on the fact that the krope learning procedure can potentially mitigate the divergence of fqe’s learning procedure depending on which pairs of transitions are sampled.
Takeaway #$\#3$: Understanding krope Divergence Depending on which pairs of transitions are sampled, krope can potentially mitigate or worsen the divergence of fqe.
C.3.2 Learning Graphs for Offline Policy Evaluation on dmc and d4rl Datasets
@@
On the d4rl dataset, we reach the similar conclusions: krope is effective in producing stable and accurate ope estimates. However, in $3/9$ instances, krope does diverge. This divergence is likely related to the discussion in Section 5 and Appendix C.3.1. Recall that krope is a semi-gradient method, which does not optimize any objective function and is susceptible to divergence (Feng et al., 2019; Sutton & Barto, 2018). So while krope representations stabilize value function learning, krope’s learning algorithm may diverge and not converge to krope representations. However, regardless of this result, krope does improve the stability and accuracy of fqe in all cases.
C.3.3 Stability-related Analysis on Custom Datasets
@@
Another metric that is associated with stability is Bellman completeness (bc) (Chang et al., 2022; Wang et al., 2021a). We find that krope is approximately Bellman complete even though it does not explicitly optimize for it; this finding aligns with our Theorem 2. While bc is difficult to approximate, we can minimize the proxy metric introduced given in Equation (10) (Chang et al., 2022):

### v4

@@ C.2 Baselines
It is a bisimulation-based metric learning algorithm that is off-policy evaluation variant of mico (Castro et al., 2021; Pavse & Hanna, 2023a). It learns representations directly such that the $\pi$-bisimulation metric is satisfied and makes no assumptions on the transition dynamics. Its additional learning rate is the output dimension of $\phi$. Furthermore, we provide additional details on our experimental setup is different from that in (Pavse & Hanna, 2023a). The main differences are:
  1. 1.
    The original rope used rope as a pre-training step, fixed the representations, and then fed them into fqe for ope. In our case, we also pre-train the representations but with fqe as a representation learning algorithm (for value predictive representations (Lehnert & Littman, 2020)) along with other representation learning algorithms as auxiliary tasks. The fixed learned representations are then fed into lspe for ope.
  2. 2.
    The original rope encoder architecture had a tanh activation function on the output layer, which we effectively serves a clipping mechanism to the features, which is similar to how public implementations of fqe clip the return to avoid divergence [^4] https://github.com/google-research/google-research/blob/master/policy_eval/q_fitter.py.
We opted to use our alternative setup for two reasons: 1) By using lspe as the ope algorithm, instead of fqe, we can precisely quantify the stability properties of the representations in terms of the spectral radius (Theorem 1), which is harder to do when using fqe as the ope algorithm; 2) while a valid architectural choice, for this current work, we viewed the use of the tanh function as obfuscating the true stability properties of the representations and so opted to avoid it. More practically, it is reasonable to use the tanh as part of the architecture.
@@
The dr3 regularizer minimizes the total feature co-adaptation by adding the term $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$ as an auxiliary task to the main fqe loss (Kumar et al., 2022). (Ma et al., 2024) introduced an improvement to this auxiliary loss by suggesting that the absolute value of the feature co-adaptation be minimized, i.e., $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}|\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})|$. We use $\alpha=0.1$ as its auxiliary task weight. Absolute dr3’s only hyperparameters are the auxiliary task weight $\alpha$ and the $\phi$ output dimension.
@@ C.3 Additional Results
C.3.1 Learning Graphs for Offline Policy Evaluation on dmc and d4rl Datasets
@@
On the d4rl dataset, we reach the similar conclusions: krope is effective in producing stable and accurate ope estimates. However, in $3/9$ instances, krope does diverge. This divergence is likely related to the discussion in Section 5 and Section 4.4. Recall that krope is a semi-gradient method, which does not optimize any objective function and is susceptible to divergence (Feng et al., 2019; Sutton & Barto, 2018). So while krope representations stabilize value function learning, krope’s learning algorithm may diverge and not converge to krope representations. However, regardless of this result, krope does improve the stability and accuracy of fqe in all cases.
C.3.2 Stability-related Analysis on Custom Datasets
@@
Another metric that is associated with stability is Bellman completeness (bc) (Chang et al., 2022; Wang et al., 2021). We find that krope is approximately Bellman complete even though it does not explicitly optimize for it; this finding aligns with our Theorem 2. While bc is difficult to approximate, we can minimize the proxy metric introduced given in Equation (10) (Chang et al., 2022):